OpenStudy (anonymous):

A 0.200 kg block of a substance requires 3.59 kJ of heat to raise its temperature by 20.0 K. What is the specific heat of the substance? A. 2,020 J/(kg * K) B. 383 J/(kg*K) C. 130 J/(kg * K) D. 897 J/(kg*K)

OpenStudy (michele_laino):

the requested specific heat c is: \[\large c = \frac{1}{m}\frac{Q}{{\Delta t}} = \frac{1}{{0.2}}\frac{{3.59}}{{20}} = ...\left( {Kjoules/Kg \times K} \right)\]

OpenStudy (anonymous):

we get 0.8975? and converting we get choice D as our answer?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thansk!!

OpenStudy (michele_laino):

:)